# Searching for (sharp) thresholds in random structures: where are we now?

### v1

@@ 1.2.3 Statistical physics
Much of the language and intuition around thresholds and sharp thresholds comes from physical systems that exhibit sudden changes in qualitative behavior as some parameter is changed in a small way. This is the phenomenon of a phase transition. The phase transitions of water from liquid to solid as temperature drops below $32^{\circ}\text{ F}$ or from liquid to gas as temperature rises past $100^{\circ}\text{ F}$ are familiar to everyone.
@@ 2 General questions, notions, and phenomena
The following result of Bollobas and Thomason justifies the abstract study of thresholds in monotone properties and the specific definition of $p_{c}(\mathcal{F})$ above.
Theorem 1 (Bollobas–Thomason [27]).
@@ 3.3 Applications
[…]
@@
[…]
@@ 4 Random $k$-SAT and the satisfiability conjecture
This is one motivation for the study of average-case complexity: the computational tractability or intractability of random instances of computational problems. Average-case complexity is a huge topic with many fascinating facets and many mysteries (see e.g. [22, 126]). Here we will deal will a small slice of the topic.
@@
The $k$-SAT problem is to find an assignment of True and False to the boolean variables so that the given $k$-CNF formula evaluates to true; if such an assignment exists the formula is satisfiable and unsatisfiable otherwise.
An empirical observation about algorithms brought the study of random computational problems to the attention of computer scientists, statistical physicists, and probabilists. In 1996, Mitchell, Selman, and Levesque [120] (working on problems in Artificial Intelligence and computational deductive reasoning) generated uniformly random 3-SAT instances on $n$ variables with $m$ constraints, for fairly large $n$ and different values of $m$. They ran standard heuristic SAT solving and SAT refutation algorithms and observed the following: the running time required to find a solution (or find a proof that none existed) showed a sharp peak (as a function of $m$) right around the point at which an estimate of the probability of such a random instance being satisfiable made a sharp decrease from near $1$ to near 00.
@@
How can one locate the satisfiability threshold in a random CSP? As with many of the examples above, a starting point is to try the first- and second-moment methods on the random variable $Z$ that counts the number of solutions. If $p$ is large enough that $\mathbb{E}Z\to 0$, we know $p_{c}\leq p$; conversely if $p$ is such that $\mathbb{E}Z\to\infty$ and, say, $\mathbb{E}(Z^{2})\leq C(\mathbb{E}Z)^{2}$, then we know $p_{c}\geq p$.
@@ 4.1 Friedgut’s Theorem
Though motivated by the Satisfiability Conjecture, Friedgut proved much more in [54]: he gave a very general characterization of what monotone properties of random graphs can have a coarse threshold: those that are well approximated by the property of containing some bounded-size subgraph from a bounded-size list of subgraphs. All other properties have a sharp threshold (in the non-uniform sense of Theorem 12).
@@ 4.2 The cavity method and the structure of solution spaces
Statistical physicists soon turned their attention to random $k$-SAT and related models using tools and intuition from the study of spin glasses [99, 101, 100, 97, 92, 102]. The replica and cavity methods are powerful analytic tools for analyzing disordered (random) systems, based on assumptions that to a large extent have not been proved mathematically. Under these assumptions, these methods are capable of making very detailed predictions for thresholds and phase transitions in a broad class of random structures, Gibbs measures on random graphs.
@@
In terms of identifying the satisfiability threshold, condensation is important because it presents a fundamental obstacle to applying the second-moment method to the random variable $Z$ counting the number of solutions: no amount of conditioning can reduce the variance sufficiently and so another approach is required (see the discussion in [39, 38, 37, 20]).
@@ 4.3 Proof of the satisfiability conjecture for large $k$
The innovation of Coja-Oghlan and Panagiotou in [39] was to design a random variable $Z_{\beta}$ counting the number of solutions with $\beta n$ ‘blocked’ variables that is a stand-in for counting the number of clusters of a given (exponential) size. The same authors apply this idea to random $k$-SAT in [37], additionally conditioning on typical vertex degrees. This kind of random variable was also used in [44, 43] to determine thresholds and limiting values of optimization problems in random regular models (in which each vertex or variable has the same edge or constraint degree).
@@ 4.5 Scaling windows
The nature of the solution space in this model raises a lot of questions: if almost all solutions are isolated should the search problem be hard at all densities? In fact there are efficient search algorithms at low densities [87] even when almost all solutions are isolated; these algorithms in fact find solutions lying in very rare clusters (with maximum possible diameter) [2], as predicted by physicists working on machine learning problems [19]. This indicates that the cartoon in Figure 6 is not really relevant to algorithms: the cartoon depicts properties of typical solutions, while efficient algorithms may indeed find rare solutions. The structural properties like the OGP that apply to all solutions (or tuples of solutions) are more algorithmically relevant.
@@ 5 Statistical inference and the stochastic block model
Though we have left the setting of monotone properties of the hypercube under $\mu_{p}$, we have retained some of the essential features: this model has independent edges and an important monotonicity property. If we fix the ratio $\frac{p_{\text{in}}}{p_{\text{out}}}=:\alpha$ and set $p_{\text{in}}=p\alpha$ and $p_{\text{out}}=p$ then the weak recovery problem only becomes easier as $p$ increases: we have more data (observed edges) from which to deduce the signal (the underlying partition).
@@ 5.1 Sharp thresholds for inference
There is an information theoretic threshold at $d_{\text{inf}}$ if
@@
For the case $q=2,3$ they conjectured that for any $\alpha$, $d_{\text{inf}}=d_{\text{alg}}$, but for $q\geq 5$ they conjectured a gap: $d_{\text{inf}}<d_{\text{alg}}$ (now known as a statistical–computational gap, and the subject of great recent interest in statistics and computer science). The $q=2$ conjecture was proved in celebrated works of Mossel–Neeman–Sly [105, 106] and Massoulié [95]
The positive side of their general conjecture about $d_{\text{alg}}$, that weak recovery is possible above the ‘Kestum–Stigum threshold’, was proved by Abbe and Sandon [5]. The information–theoretic threshold conjecture in the disassortative case ($\alpha<1$) was proved in [36] by implementing a rigorous version of the cavity method, but the assortative case remains open in general. The surveys [104, 1] contain detailed accounts of further developments.
@@ 6 Conclusions and questions
Just as in $G(n,p)$, a general study of monotone properties is possible in $G_{d}(n,p)$. In particular, let $\mathcal{F}$ be a non-trivial monotone property (where as always monotone pertains to adding edges). Then again the probability that $\mathcal{F}$ holds in $G_{d}(n,p)$ is a strictly increasing function of $p$, and so $p_{c}(\mathcal{F})$ can be uniquely defined, and all of the same questions from Section 1.1 can be asked in this setting too.
Perhaps surprisingly then, the analogue of Theorem 1, the Bollobas-Thomason theorem, that lays the foundation for the general study of thresholds of monotone properties in random graphs, is not known in general for random geometric graphs. In the special case of $d=1$ (points on a circle), McColm proved that every monotone property has a threshold [96]. Moreover, Goel, Rai, and Krishnamachari [66] proved that the scaling window of every monotone property is bounded by a function (depending on $d$) that vanishes as $n\to\infty$, the analogue of the result of Friedgut and Kalai [56] for $G(n,p)$. The general statement, however, remains open.

### v2

@@ 1.2.3 Statistical physics
Much of the language and intuition around thresholds and sharp thresholds comes from physical systems that exhibit sudden changes in qualitative behavior as some parameter is changed in a small way. This is the phenomenon of a phase transition. The phase transitions of water from liquid to solid as temperature drops below $32^{\circ}\text{ F}$ or from liquid to gas as temperature rises past $212^{\circ}\text{ F}$ are familiar to everyone.
@@ 2 General questions, notions, and phenomena
The following result of Bollobás and Thomason justifies the abstract study of thresholds in monotone properties and the specific definition of $p_{c}(\mathcal{F})$ above.
Theorem 1 (Bollobás–Thomason [29]).
@@ 3.3 Applications
The property $\mathcal{F}_{\text{Latin}}$ has a sharp threshold at $p=\frac{\log n}{n}$
@@
$p_{c}(\mathcal{F}_{\text{Latin}})=\Theta\left(\frac{\log n}{n}\right)\,.$
@@ 4 Random $k$-SAT and the satisfiability conjecture
This is one motivation for the study of average-case complexity: the computational tractability or intractability of random instances of computational problems. Average-case complexity is a huge topic with many fascinating facets and many mysteries (see e.g. [24, 135]). Here we will deal with a small slice of the topic.
@@
The $k$-SAT decision problem is to determine if there exists an assignment of True and False to the boolean variables so that the given $k$-CNF formula evaluates to true; if such an assignment exists the formula is satisfiable and unsatisfiable otherwise.
An empirical observation about algorithms brought the study of random computational problems to the attention of computer scientists, statistical physicists, and probabilists. In 1992, Mitchell, Selman, and Levesque [107] (working on problems in Artificial Intelligence and computational deductive reasoning) generated uniformly random 3-SAT instances on $n$ variables with $m$ constraints, for fairly large $n$ and different values of $m$. They ran standard heuristic SAT solving and SAT refutation algorithms and observed the following: the running time required to find a solution (or find a proof that none existed) showed a sharp peak (as a function of $m$) right around the point at which an estimate of the probability of such a random instance being satisfiable made a sharp decrease from near $1$ to near 00.
@@
How can one locate the satisfiability threshold in a random CSP? As with many of the examples above, a starting point is to try the first- and second-moment methods on the random variable $Z$ that counts the number of solutions. If $p$ is large enough that $\mathbb{E}Z\to 0$, we know $p_{c}\leq p$; conversely if $p$ is such that $\mathbb{E}Z\to\infty$ and, say, $\mathbb{E}(Z^{2})\leq C(\mathbb{E}Z)^{2}$, then (via the Paley–Zygmund Inequality and Friedgut’s Theorem, see Sections 4.1 and 4.2) we know $p_{c}\geq p$.
@@ 4.1 Friedgut’s Theorem
Though motivated by the Satisfiability Conjecture, Friedgut proved much more in [57]: he gave a very general characterization of which monotone properties of random graphs can have a coarse threshold: those that are well approximated by the property of containing some bounded-size subgraph from a bounded-size list of subgraphs. All other properties have a sharp threshold (in the non-uniform sense of Theorem 12).
@@ 4.2 The cavity method and the structure of solution spaces
Statistical physicists soon turned their attention to random $k$-SAT and related models using tools and intuition from the study of spin glasses [105, 108, 106, 103, 96, 109]. The replica and cavity methods they developed are powerful analytic tools for analyzing disordered (random) systems, based on assumptions that to a large extent have not been proved mathematically. Both are methods for computing $\mathbb{E}\log Z$, the expectation of a log partition function, the normalizing constant of a Gibbs measure [^2] Here the setting is always at positive temperature so $Z>0$ and taking a logarithm is justified; models with hard constraints like random $k$-SAT or $q$-coloring are analyzed by taking a limit as the temperature tends to 00.. The replica method is based on the identity $\log Z=\lim_{n\to 0}\frac{Z^{n}-1}{n}$, and rather magically using computations for large integer $n$ to deduce the value of the limit. See [134, Chapter 1.13] for a mathematical introduction. The cavity method is based on a coupling of random graphs or hypergraphs on $n$ vertices and on $n+1$ vertices, and understanding the expected change in the partition function, $\log Z_{n+1}-\log Z_{n}$, as a function of the Gibbs measure and its correlation properties. See [104, Chapter 19] for an introduction to the method. Under strong assumptions, these two methods are capable of making very detailed predictions for thresholds and phase transitions in a broad class of random structures, Gibbs measures on random graphs.
@@
In terms of identifying the satisfiability threshold, condensation is important because it presents a fundamental obstacle to applying the second-moment method to the random variable $Z$ counting the number of solutions. We briefly describe the connection at a high level. Let $d$ denote the average vertex or variable degree of a random CSP. The exponential order of the first moment, $f_{1}(d):=\lim_{n\to\infty}\frac{1}{n}\log\mathbb{E}Z$, is a decreasing function of $d$ and the first-moment threshold $d_{\mathrm{first}}$ is the point at which it hits 00. For $d>d_{\mathrm{first}}$ whp there are no solutions by Markov’s Inequality; for $d<d_{\mathrm{first}}$ the expected number of solutions is exponentially large. To show that there is at least one solution with probability bounded away from 00, we would like to apply the second-moment method via the Paley–Zygmund Inequality: $\Pr(Z>0)\geq\frac{(\mathbb{E}Z)^{2}}{\mathbb{E}(Z^{2})}$. Writing $\mathbb{E}(Z^{2})=\sum_{\sigma,\tau}\Pr[\sigma,\tau\text{ both solutions}]$ one sees that on an exponential scale, the second moment is at most the square of the first moment if and only if the dominant contribution to $\mathbb{E}(Z^{2})$ comes from pairs $\sigma,\tau$ of uncorrelated vectors. Past the condensation threshold, pairs of solutions with non-trivial correlation are significant and so the second moment should be expected to be exponentially larger than the square of the first moment. See the discussion in [10, 41, 40, 39, 22] for much more detail. To prove a random CSP is whp satisfiable beyond the condensation threshold requires a significantly different approach.
@@ 4.3 Proof of the satisfiability conjecture for large $k$
The innovation of Coja-Oghlan and Panagiotou in [41] was to design (for the closely related random $k$-NAE SAT model333‘not-all-equal’ SAT: an assignment is satisfying if every clause has at least one true and at least one false literal.) a random variable $Z_{\beta}$ counting the number of solutions with $\beta n$ ‘blocked’ variables that is a stand-in for counting the number of clusters of a given (exponential) size. The same authors apply this idea to random $k$-SAT in [39], additionally conditioning on typical vertex degrees. This kind of random variable was also used in [46, 45] to determine thresholds and limiting values of optimization problems in random regular models (in which each vertex or variable has the same edge or constraint degree).
@@ 4.5 Scaling windows
The nature of the solution space in this model raises a lot of questions: if almost all solutions are isolated should the search problem be hard at all densities? In fact there are efficient search algorithms at low densities [91, 2] even when almost all solutions are isolated; these algorithms in fact find solutions lying in very rare clusters (with maximum possible diameter) [2], as predicted by physicists working on machine learning problems [20]. These algorithms are ‘multiscale majority algorithms’ in which entries of a potential solution vector $X$ are set round by round to optimize a greedy objective. The symmetric binary perceptron is a random model of the combinatorial discrepancy problem, and so constructive discrepancy minimization algorithms also provide provably efficient search algorithms at low constraint densities [99, 125, 50, 21]. These algorithms can also be adapted for the original (non-symmetric) perceptron model [97]. The success of these algorithms indicates that the cartoon in Figure 6 is not really relevant to algorithms: the cartoon depicts properties of typical solutions, while efficient algorithms may indeed find rare solutions. The structural properties like the OGP that apply to all solutions (or tuples of solutions) are more algorithmically relevant.
@@ 5 Statistical inference and the stochastic block model
Though we have left the setting of monotone properties of the hypercube under $\mu_{p}$, we have retained some of the essential features: this model has independent edges and an important monotonicity property. If we fix the ratio $\frac{p_{\text{in}}}{p_{\text{out}}}=:\alpha$ and set $p_{\text{in}}=p\alpha$ and $p_{\text{out}}=p$ then the weak recovery problem only becomes easier as $p$ increases: we have more data (observed edges) from which to deduce the signal (the underlying partition). That is, if $p>p^{\prime}$ we can reduce the model with parameter $p$ to that with parameter $p^{\prime}$ by independently deleting each edge with probability $p^{\prime}/p$.
@@ 5.1 Sharp thresholds for inference
Fix a number of parts $q\geq 2$ and a ratio $\alpha>0$ of $p_{\text{in}}/p_{\text{out}}$. There is an information theoretic threshold at $d_{\text{inf}}$ if
@@
For the case $q=2,3$ they conjectured that for any $\alpha$, $d_{\text{inf}}=d_{\text{alg}}$, but for $q\geq 5$ they conjectured a gap: $d_{\text{inf}}<d_{\text{alg}}$ (now known as a statistical–computational gap, and the subject of great recent interest in statistics and computer science). In general the conjectured value $d_{\text{alg}}$ is explicit (it is the ‘Kestum–Stigum threshold’ [89]), while the conjectured value of $d_{\text{inf}}$ is given implicitly as the solution of a complicated optimization problem over probability measures. The $q=2$ conjectures were proved in celebrated works of Mossel–Neeman–Sly [112, 113] and Massoulié [101].
For $q\geq 3$, partial progress has been made. Abbe and Sandon [5] proved that for any $q$ and any $\alpha$, weak recovery is possible above the Kestum–Stigum threshold, i.e. the positive part of the conjecture on the algorithmic threshold. For $q\geq 3$ and $\alpha<1$ (the disassortative case), the conjecture on the value of the information–theoretic threshold (as the solution of the complicated optimization problem) was proved in [38] by implementing a rigorous version of the cavity method. Recently, it was proved in [116] that for $q=3$ (and for $q=4$ and $\alpha$ sufficiently close to $1$), $d_{\text{inf}}=d_{\text{alg}}$.
@@ 6 Conclusions and questions
Just as in $G(n,p)$, a general study of monotone properties is possible in $G_{d}(n,p)$. In particular, let $\mathcal{F}$ be a non-trivial monotone property (where, as always, monotone pertains to adding edges). Then again the probability that $\mathcal{F}$ holds in $G_{d}(n,p)$ is a strictly increasing function of $p$, and so $p_{c}(\mathcal{F})$ can be uniquely defined, and all of the same questions from Section 1.1 can be asked in this setting too.
Perhaps surprisingly then, the analogue of Theorem 1, the Bollobás-Thomason theorem, that lays the foundation for the general study of thresholds of monotone properties in random graphs, is not known in general for random geometric graphs. In the special case of $d=1$ (points on a circle), McColm proved that every monotone property has a threshold [102]. Moreover, Goel, Rai, and Krishnamachari [69] proved that the scaling window of every monotone property is bounded by a function (depending on $d$) that vanishes as $n\to\infty$, the analogue of the result of Friedgut and Kalai [59] for $G(n,p)$. The general statement, however, remains open.